he needs to make 60 liters of a 36% acid solution to test a new product. His supplier only ships a 10% and 40% solution. he decides to make the 36% solution by mixing the 10% solution with the 40% solution. how much of the 10% solution will he need to use
Found 2 solutions by Alan3354, greenestamps:
Answer by Alan3354(69443)

he needs to make 60 liters of a 36% acid solution to test a new product. His supplier only ships a 10% and 40% solution. he decides to make the 36% solution by mixing the 10% solution with the 40% solution. how much of the 10% solution will he need to use
====================================
Similar problem:
-----
Birdseed costs ​$0.56 a pound and sunflower seeds cost ​$0.86 a pound. Angela​ Leinenbachs' pet store wishes to make a 40​-pound mixture of birdseed and sunflower seeds that sells for ​$0.75 per a pound. How many pounds of each type of seed should she​ use?
==================
b + s = 40 --- total mix
56b + 86s = 40*75 ---- cost in cents

Answer by greenestamps(13200)



A quick and easy non-algebraic way to solve 2-part mixture problems like this:

(1) 36% is 26/30 of the way from 10% to 40% (from 10 to 40 is 30; from 10 to 36 is 26...)
(2) That means 26/30 of the mixture has to be the 40% solution

Since he is making 60 liters of the mixture, he needs 26/30 of 60 liters, or 52 liters, of the 40% solution; the other 8 liters are the 10% solution.

CHECK:
.40(52)+.10(8) = 20.8+0.8 = 21.6
.36(60) = 21.6